# Mixed Berndt-Type Integrals and Generalized Barnes Multiple Zeta Functions

### v2

@@
Abstract. In this paper, we define and study four families of Berndt-type integrals, called mixed Berndt-type integrals, which contains (hyperbolic) sine and cosine functions in the integrand function. By contour integration, these integrals are first converted to some hyperbolic (infinite) sums of Ramanujan type, all of which can be calculated in closed forms by comparing both the Fourier series expansions and the Maclaurin series expansions of a few Jacobi elliptic functions. These sums can be expressed as rational polynomials of $\Gamma(1/4)$ and $\pi^{-1}$ which give rise to the closed formulas of the mixed Berndt-type integrals we are interested in. Moreover, we also present some interesting consequences and illustrative examples. Further, we define a generalized Barnes multiple zeta function, and find a classic integral representation of the generalized Barnes multiple zeta function. Furthermore, we give an alternative evaluation of the mixed Berndt-type integrals in terms of the generalized Barnes multiple zeta function. Finally, we obtain some direct evaluations of rational linear combinations of the generalized Barnes multiple zeta function.
@@ 1 Introduction
where $s\geq 1$ and $m\geq 1$ if the denominator has “$+$’ sign, and $s\geq 2m+1$ otherwise. This kind of improper integral can be traced back to the esteemed mathematician S. Ramanujan. First as a problem submitted by Ramanujan [14, pp. 325-326] over a century ago to the Indian Journal of Pure and Applied Mathematics:
@@
Here $n$ is any odd positive integer. Four year after submission, Wilkinson [19] provided a proof. About 80 years later, Ismail and Valent [10] found the mystery integral (Ismail-type integrals)
@@
where $K\equiv K(x)$ denotes the complete elliptic integral of the first kind and $K^{\prime}=K(1-x)$, see Section 2. The integrand in (1.2) is the weight function for a sequence of polynomials arising in the Hamburger moment problem studied by Ismail and Valent. The systematic study of Berndt-type integrals began with the works of Kuznetsov [12] and Berndt [6] about ten years ago. Some recent results on infinite series involving Berndt-type integrals may be found in the works of [7, 13, 16, 17, 21, 22, 24] and the references therein. Kuznetsov [12] used the contour integrations and theta functions to prove a more general result of Ismail-type integrals
@@
where ${\rm cn}(u,k):=\sqrt{1-{\rm sn}^{2}(u,k)}$ and ${\rm dn}(u,k):=\sqrt{1-k^{2}{\rm sn}^{2}(u,k)}$. Berndt [6] provided a direct evaluation of (1.2) and other similar integrals of this type by using Cauchy’s residue theory, the Fourier series expansions, and the Maclaurin series expansions of a few Jacobi elliptic functions. In particular, using Berndt’s results [6] and Xu-Zhao’s result [22], it is easy to see that the Berndt-type integrals of order one have the following structures
@@
Moreover, Berndt [6, Thm. 6.1] also considered some integrals contain $\sinh x\pm i\sin x$ in their integrands instead of $\cosh x\pm\cos x$. Further, using the contour integrals and the Fourier series expansion and Maclaurin series expansion of a certain Jacobi elliptic function, Rui-Xu-Yang-Zhao obtained a structural result on the original Berndt type integral involving (hyperbolic) sine functions (see [16, Thm. 3.4])
@@
and prove the following result involving another family of Berndt-type integrals (see [16, Cor. 6.5]):
@@
Recently, by extending an argument as used in the proof of the main theorems of [6], Xu and Zhao [21] proved that the following explicit evaluations of two families of Berndt-type integrals of order two (cf. [21, Thm. 1.3]):
@@
Further, Rui, Xu and Zhao [17, Thm. 1.1] established explicit evaluations of two families of Berndt-type integrals of order three by special values of the Gamma function. More generally, let $X=\Gamma^{4}(1/4)$ and $Y=\pi^{-1}$, and $\mathbb{Q}[x,y]$ denotes all rational linear combinations of products of $x$ and $y$ (the polynomial ring generated by $x$ and $y$ with rational coefficients). Xu and Zhao [21, 22] proved two structural theorems on the general Berndt-type integrals with the denominator having arbitrary positive degrees. Specifically, they shown that for all integers $m\geq 1$ and $p\geq[m/2]$, the Berndt-type integrals satisfy
@@
where the degrees of $X$ have the same parity as that of $m$ and are between $22p-m+$ 2 and $2p+m$, inclusive, while the degrees of $Y$ are between $22p-m+$ 2 and $22p+3m-$ 2, inclusive. Similarly, they shown that
@@
for all $p\geq m+1\geq 1$ and that the degrees of $X$ in each term are always even and between $2p-2m$ and $2p+2m$, while the degrees of $Y$ are between $2p-2m$ and $2p+6m$. In the even order case, they shown that
@@
for all integers $p\geq m\geq 1$ and the powers of $X$ are between $2p+2-2m$ and $2p+2m$, while the degrees of $Y$ are between $2p+2-2m$ and $22p+6m-$ 2. Very recently, it is very surprising that Bradshaw and Vignat [7, Cor. 2] discovered a close connection between Berndt-type integrals and Barnes zeta functions. Moreover, Bradshaw and Vignat [7, Cor. 5] used the Kuznetsov’s method to prove the following relations between Jacobi elliptic functions and another Ismail-type integrals: for $n\in\mathbb{N}_{0}$,
@@
From the above, it can be seen that the integrand in the Berndt-type integrals studied above either only contain (hyperbolic) sine functions or (hyperbolic) cosine functions. In this paper, we consider the following infinite integrals containing (hyperbolic) sine functions and (hyperbolic) cosine functions in their integrands:
@@
We call them mixed Berndt-type integrals. We use the contour integrals and the series associated with Jacobi elliptic functions to establish two structural theorems of mixed Berndt-type integrals. We will prove the following evaluations: let $\Gamma:=\Gamma(1/4)$, for $m\in\mathbb{N}$,
@@
where $\boldsymbol{\sigma}_{N}=\left(\sigma_{1},\ldots,\sigma_{N}\right)\in\left\{\pm 1\right\}^{N}$. We find a classic integral representation:
@@
and establish some explicit relations between the mixed Berndt-type integrals and (generalized) Barnes multiple zeta functions. Finally, we establish some closed form evaluations of the combinations of several generalized Barnes multiple zeta function.
2 Some Preliminary Results
Let ${}_{2}F_{1}\left(a,b;c;x\right)$ denote the Gaussian hypergeometric function that
${{}_{2}}F_{1}(a,b;c;x)=\displaystyle\sum\limits_{n=0}^{\infty}\displaystyle\frac{(a)_{n}(b)_{n}}{(c)_{n}}\displaystyle\frac{x^{n}}{n!}\quad(a,b,c\in\mathbb{C}).$
Here $\left(a\right)_{0}:=1$ and $\left(a\right)_{n}=\displaystyle\frac{\Gamma\left(a+n\right)}{\Gamma\left(a\right)}=\prod_{j=0}^{n-1}{\left(a+j\right)}\quad(n\in\mathbb{N}),$ and $\Gamma(x)$ is the Gamma function defined by
$\Gamma\left(z\right):=\displaystyle\int_{0}^{\infty}{e^{-t}t^{z-1}\mathrm{d}t}\quad\left(\Re\left(z\right)>0\right).$
We know that ${}_{2}F_{1}\left(a,b;c;x\right)$ has numerous applications in mathematics, physics and even engineering. In particular, it can be used to evaluate some important elliptic integrals. For example, the complete elliptic integral of the first kind is defined by
$K:=K\left(k^{2}\right)=\displaystyle\int_{0}^{\pi/2}{\displaystyle\frac{\mathrm{d}\varphi}{\sqrt{1-k^{2}\sin^{2}\varphi}}=\displaystyle\frac{\pi}{2}}\,\,_{2}F_{1}\left(\displaystyle\frac{1}{2},\displaystyle\frac{1}{2};1;k^{2}\right),$
where $x=k^{2}\quad(0<k<1)$ is the modulus of $K$. Let $k^{\prime}=\sqrt{1-k^{2}}$, and
$K^{\prime}:=K\left(k^{\prime 2}\right)=\displaystyle\int_{0}^{\pi/2}{\displaystyle\frac{\mathrm{d}\varphi}{\sqrt{1-k^{\prime 2}\sin^{2}\varphi}}=\displaystyle\frac{\pi}{2}}\,\,_{2}F_{1}\left(\displaystyle\frac{1}{2},\displaystyle\frac{1}{2};1;1-k^{2}\right).$
Similarly, the complete elliptic integral of the second kind is denoted by (Whittaker and Watson [18])
$E:=E(k^{2}):=\displaystyle\int_{0}^{\pi/2}\sqrt{1-k^{2}\sin^{2}\varphi}\mathrm{d}\varphi=\displaystyle\frac{\pi}{2}{{}_{2}}F_{1}\left(-\displaystyle\frac{1}{2},\displaystyle\frac{1}{2};1;k^{2}\right).$
@@ 2 Some Preliminary Results
Then
$\frac{\mathrm{d}^{n}z}{\mathrm{d}x^{n}}=\displaystyle\frac{\left(1/2\right)_{n}^{2}}{n!}\,_{2}F_{1}\left(\displaystyle\frac{1}{2}+n,\displaystyle\frac{1}{2}+n,1+n;x\right).$
Applying the identity ([2])
$\,{}_{2}F_{1}\left(a,b,\displaystyle\frac{a+b+1}{2};\displaystyle\frac{1}{2}\right)=\displaystyle\frac{\Gamma\left(\displaystyle\frac{1}{2}\right)\Gamma\left(\displaystyle\frac{a+b+1}{2}\right)}{\Gamma\left(\displaystyle\frac{a+1}{2}\right)\Gamma\left(\displaystyle\frac{b+1}{2}\right)},$
we can calculate that
$\frac{d^{n}z}{dx^{n}}\mid_{x=\tfrac{1}{2}}=\displaystyle\frac{\left(1/2\right)_{n}^{2}\sqrt{\pi}}{\Gamma^{2}\left(\displaystyle\frac{n}{2}+\displaystyle\frac{3}{4}\right)}.$
3 Some Definitions and Lemmas
@@
([21, Lem. 7.3.]) For any integer $m\geq 0$ there is a polynomial $g_{m}(X,Y,Z)\in\mathbb{Z}[X,Y,Z]$ such that
$\displaystyle\frac{\mathrm{d}^{m}}{\mathrm{d}u^{m}}\mathrm{sn}\left(u\right)=g_{m}\left(\sin\varphi,\cos\varphi,k^{2}\right)\cdot\begin{cases}\left(1-k^{2}\sin^{2}\varphi\right)^{1/2},if\,\,m\,\,is\,\,odd;\\ 1,if\,\,m\,\,is\,\,even,\\ \end{cases}$ (3.15)
@@
([23, Thm.2.4]) Let $x,y,z$ and $z^{\prime}$ satisfy (2.14). Given the formula $\Omega(x,e^{-y},z,z^{\prime})=0$, we have the transformation formulas
@@
$\displaystyle\frac{\left(-1\right)^{n-1}}{\cos\left(\displaystyle\frac{1\pm i}{2}z\right)}\overset{z\rightarrow\left(1-i\right)\tilde{n}\pi}{=}\displaystyle\frac{1\mp i}{z-\left(1\mp i\right)\tilde{n}\pi}-2\displaystyle\sum_{k=1}^{\infty}{\displaystyle\frac{\overline{\zeta}\left(2k\right)}{\pi^{2k}}\left(\displaystyle\frac{1\pm i}{2}\right)^{2k-1}\left(z-\left(1\mp i\right)\tilde{n}\pi\right)^{2k-1}},$ (3.20)
where $\tilde{n}:=n-\displaystyle\frac{1}{2}$. Here $\zeta(s)$ and $\overline{\zeta}(s)$ stand for the Riemann zeta function and alternating Riemann zeta function, respectively, which are defined by
$\zeta\left(s\right):=\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{1}{n^{s}}}\quad(\Re(s)>1)\quad\text{and}\quad\bar{\zeta}\left(s\right):=\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{\left(-1\right)^{n-1}}{n^{s}}}\quad\left(\Re\left(s\right)>0\right).$
When $s=2m$ is an even, Euler proved the famous formula
$\zeta\left(2m\right)=\displaystyle\frac{\left(-1\right)^{m-1}B_{2m}\left(2\pi\right)^{2m}}{2\left(2m\right)!},$
where $B_{2m}\in\mathbb{Q}$ are Bernoulli numbers defined by the generating function
$\frac{x}{e^{x}-1}=\displaystyle\sum_{n=0}^{\infty}{\displaystyle\frac{B_{n}}{n!}x^{n}.}$
@@
([23, Lem. 1]) Let $p$ be a positive integer and $\alpha,\beta$ be real numbers such as $\alpha\beta=\pi^{2}$
@@ 4 Mixed Berndt-Type Integrals via Hyperbolic Series
In this section, we will establish some explicit relations between the mixed Berndt-type integrals and some hyperbolic summations by using the contour integrations.
@@
where $C_{R}$ denotes the quarter- circular contour consisting of the interval $[0,R]$, the qiarter-circle $\Gamma_{R}$ with $|z|=R,0\leq argz\leq\pi/2$, and $[iR,0]$, please see the following figure:
Obviously, there exist poles in case that
$\left(\sinh z+i\sin z\right)\left(\cosh x+\cos x\right)=2i\cos^{2}\left\{\displaystyle\frac{1+i}{2}z\right\}\sin\left\{\left(1-i\right)z\right\}=0.$
The only lying inside $C_{R}$ are $z_{n}=n(1+i)\pi/2\ (n\geq 1,\ |z_{n}|<R)$. Using Lemma 3.3, the residue $\text{Res}[F_{1}(z),z_{n}]$ at each such pole is give by
@@
As $R\rightarrow\infty$, it is to show that
@@
Thus, combining related identities, we deduce the desired result. ∎
@@
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{a}\left(\sinh x-i\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}+$ $\left(-1\right)^{a}i^{a}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{a}\left(\sin x-i\sinh x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$
@@
where $C_{R}$ denotes the quarter-circular contour consisting of the interval $[0,R]$, the qiarter-circle $\Gamma_{R}$ with $|z|=R\ (-\pi/2\leq argz\leq 0)$, and $[-iR,0]$, please see the following figure.
Obviously, there exist poles in case that
$\left(\sinh z+i\sin z\right)\left(\cosh z-\cos z\right)=-2i\sin^{2}\left\{\displaystyle\frac{1-i}{2}z\right\}\sin\left\{\left(1+i\right)z\right\}=0.$
The only lying inside $C_{R}$ are $z_{n}=n(1-i)\pi/2\ (n\geq 1,\ |z_{n}|<R)$. Using Lemma 3.3, the residue $\text{Res}[F_{2}(z),z_{n}]$ at each such pole is give by
$\mathrm{Res}\left[F_{2}\left(z\right),z=\displaystyle\frac{1-i}{2}n\pi\right]=\displaystyle\frac{\left(-1\right)^{n}\left(\tfrac{1-i}{2}n\pi\right)^{a}}{2i\left(1+i\right)\sinh^{2}\left(\left(\tfrac{1-i}{2}\right)^{2}n\pi\right)}$
$=\displaystyle\frac{\left(-1\right)^{n}\left(\tfrac{1-i}{2}n\pi\right)^{a}}{2i\left(1+i\right)\sinh^{2}\left(\tfrac{n\pi}{2}\right)}=\left(\displaystyle\frac{1-i}{2}\right)^{a+1}\displaystyle\frac{\pi^{a}}{2i}\displaystyle\frac{\left(-1\right)^{n}n^{a}}{\sinh^{2}\left(\displaystyle\frac{n\pi}{2}\right)}.$ (4.29)
It is to show that, letting $R\rightarrow\infty$,
@@
$2\pi i\displaystyle\sum_{n=1}^{\infty}{\mathrm{Res}\left[F_{2}\left(z\right),z_{n}\right]}=\displaystyle\lim_{R\rightarrow\infty}\displaystyle\int_{\mathrm{C}_{R}}{\displaystyle\frac{z^{a}\mathrm{d}z}{\left(\sinh z+i\sin z\right)\left(\cosh z-\cos z\right)}}$
$=\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{a}\mathrm{d}x}{\left[\sinh x+i\sin x\right]\left[\cosh x-\cos x\right]}}-i\displaystyle\int_{0}^{\infty}{\displaystyle\frac{\left(-ix\right)^{a}\mathrm{d}x}{\left[\sinh\left(ix\right)+i\sin\left(ix\right)\right]\left[\cosh\left(ix\right)-\cos\left(ix\right)\right]}}$
$=\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{a}\left(\sinh x-i\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}+\displaystyle\int_{0}^{\infty}{\displaystyle\frac{(-i)^{a}x^{a}\mathrm{d}x}{\left[\sin x+i\sinh x\right]\left[\cosh x-\cos x\right]}}.$
Thus, the proof is completed. ∎
@@
where $C_{R}$ denotes the same positively oriented quarter-circular contour that we used in the proof of Theorem 4.2. The proof is completely similar as the proof of Theorem 4.2. So we leave it to the interested reader. ∎
@@
where $C_{R}$ denotes the same positively oriented quarter-circular contour that we used in the proof of Theorem 4.1. The proof is completely similar as the proof of Theorem 4.1. We leave the details to the interested reader. ∎
Setting $a=2\mu+1$ and $2\mu$ in Theorems 4.1-4.4 respectively, yield the following corollaries.

### v4

@@
Abstract. In this paper, we define and study four families of Berndt-type integrals, called mixed Berndt-type integrals, which contain (hyperbolic) sine and cosine functions in the integrand function. Using contour integration, these integrals are first converted to some hyperbolic (infinite) sums of Ramanujan type, all of which can be calculated in closed form by comparing both the Fourier series expansions and the Maclaurin series expansions of certain Jacobi elliptic functions. These sums can be expressed as rational polynomials of $\Gamma(1/4)$ and $\pi^{-1}$ which give rise to the closed formulas of the mixed Berndt-type integrals we are interested in. Moreover, we also present some interesting consequences and illustrative examples. Additionally, we define a generalized Barnes multiple zeta function, and find a classic integral representation of the generalized Barnes multiple zeta function. Furthermore, we give an alternative evaluation of the mixed Berndt-type integrals in terms of the generalized Barnes multiple zeta function. Finally, we obtain some direct evaluations of rational linear combinations of the generalized Barnes multiple zeta function.
@@ 1 Introduction
where $m\geq 1$ and $s\geq 1$ if the denominator has “++’ sign, and $s\geq 2m+1$ otherwise. This improper integral traces back to S. Ramanujan, who first submitted it as a problem over a century ago to the Indian Journal of Pure and Applied Mathematics [16, pp. 325-326]:
@@
Here, $n$ is an odd positive integer. After submission, Wilkinson [21] proved it.
About 80 years later, Ismail and Valent [10] found the mystery integral (Ismail-type integrals)
@@
where $K\equiv K(x)$ denotes the complete elliptic integral of the first kind and $K^{\prime}=K(1-x)$, see Section 2. The integrand in (1.2) is the weight function for a sequence of polynomials arising in the Hamburger moment problem studied by Ismail and Valent. The systematic study of Berndt-type integrals began with the works of Kuznetsov [12] and Berndt [6] about ten years ago. Some recent results on infinite series involving Berndt-type integrals may be found in the works of [7, 15, 18, 19, 23, 24, 26] and the references therein. Kuznetsov [12] used the contour integration and theta functions to prove a more general result of Ismail-type integrals
@@
where ${\rm cn}(u,k):=\sqrt{1-{\rm sn}^{2}(u,k)}$ and ${\rm dn}(u,k):=\sqrt{1-k^{2}{\rm sn}^{2}(u,k)}$. Berndt [6] provided a direct evaluation of (1.2) and other similar integrals of this type by using Cauchy’s residue theory, the Fourier series expansions, and the Maclaurin series expansions of certain Jacobi elliptic functions. In particular, using Berndt’s results [6] and Xu-Zhao’s result [24], it is easy to see that the Berndt-type integrals of order one have the following structures
@@
Moreover, Berndt [6, Thm. 6.1] also considered some integrals containing $\sinh x\pm i\sin x$ in their integrands instead of $\cosh x\pm\cos x$. Further, using the contour integrals and the Fourier series expansion and Maclaurin series expansion of a certain Jacobi elliptic function, Rui-Xu-Yang-Zhao obtained a structural result on the original Berndt type integral involving (hyperbolic) sine functions (see [18, Thm. 3.4])
@@
and prove the following result involving another family of Berndt-type integrals (see [18, Cor. 6.5]):
@@
Recently, by extending an argument as used in the proof of the main theorems of [6], Xu and Zhao [23] derived explicit evaluations for two families of order-two Berndt-type integrals (cf. [23, Thm. 1.3]):
@@
Further, Rui, Xu and Zhao [19, Thm. 1.1] established explicit evaluations for two families of order-three Berndt-type integrals using special values of the Gamma function. More generally, let $X=\Gamma^{4}(1/4)$ and $Y=\pi^{-1}$, and $\mathbb{Q}[x,y]$ denotes all rational linear combinations of products of $x$ and $y$ (the polynomial ring generated by $x$ and $y$ with rational coefficients). Xu and Zhao [23, 24] proved two structural theorems on the general Berndt-type integrals with the denominator having arbitrary positive degrees. Specifically, they have shown that for all integers $m\geq 1$ and $p\geq[m/2]$, the Berndt-type integrals satisfy
@@
where the degrees of $X$ have the same parity as that of $m$ and are between 2pm+22p-m+2 and $2p+m$, inclusive, while the degrees of $Y$ are between 2pm+22p-m+2 and 2p+3m22p+3m-2, inclusive. Similarly, they have shown that
@@
for all $p\geq m+1\geq 1$ and that the degrees of $X$ in each term are always even and between 2p2m2p-2m and $2p+2m$, while the degrees of $Y$ are between 2p2m2p-2m and $2p+6m$. In the even order case, they have shown that
@@
for all integers $p\geq m\geq 1$, where the powers of $X$ lie between 2p+22m2p+2-2m and $2p+2m$, while the degrees of $Y$ are between 2p+22m2p+2-2m and 2p+6m22p+6m-2. Very recently, Bradshaw and Vignat [7, Cor. 2] surprisingly discovered a close connection between Berndt-type integrals and Barnes zeta functions. Moreover, Bradshaw and Vignat [7, Cor. 5] used the Kuznetsov’s method to prove the following relation between Jacobi elliptic functions and another Ismail-type integrals: for $n\in\mathbb{N}_{0}$,
@@
From the above, it can be seen that the integrands in the Berndt-type integrals studied above either only contain (hyperbolic) sine functions or (hyperbolic) cosine functions. In this paper, we consider the following infinite integrals containing (hyperbolic) sine functions and (hyperbolic) cosine functions in their integrands:
@@
We call them mixed Berndt-type integrals. Using contour integration and series associated with Jacobi elliptic functions, we establish two structural theorems for these integrals. Specifically, we prove the following evaluations: let $\Gamma:=\Gamma(1/4)$, for $m\in\mathbb{N}$,
@@
where $\boldsymbol{\sigma}_{N}=\left(\sigma_{1},\ldots,\sigma_{N}\right)\in\left\{\pm 1\right\}^{N}$. We also derive a classic integral representation:
@@
and establish some explicit relations between mixed Berndt-type integrals and (generalized) Barnes multiple zeta functions. Finally, we establish some closed-form evaluations of the combinations of several generalized Barnes multiple zeta functions.
2 Some Definitions and Lemmas
@@
where the complete elliptic integral of the first kind is defined by
$K\equiv K(k):=\displaystyle\int_{0}^{\pi/2}{\displaystyle\frac{\mathrm{d}\varphi}{\sqrt{1-k^{2}\sin^{2}\varphi}}=\displaystyle\frac{\pi}{2}}\,\,_{2}F_{1}\left(\displaystyle\frac{1}{2},\displaystyle\frac{1}{2};1;k^{2}\right),$
additionally, let $k^{\prime}=\sqrt{1-k^{2}}$, and define the associated complete elliptic integral by
$K^{\prime}\equiv K(k^{\prime}):=\displaystyle\int_{0}^{\pi/2}{\displaystyle\frac{\mathrm{d}\varphi}{\sqrt{1-k^{\prime 2}\sin^{2}\varphi}}=\displaystyle\frac{\pi}{2}}\,\,_{2}F_{1}\left(\displaystyle\frac{1}{2},\displaystyle\frac{1}{2};1;k^{\prime 2}\right).$
@@
([23, Lem. 7.3.]) For any integer $m\geq 0$ there is a polynomial $g_{m}(X,Y,Z)\in\mathbb{Z}[X,Y,Z]$ such that
[…]
@@
([25, Thm. 2.4]) Let $x,y,z$ and $z^{\prime}$ satisfy (2.14). Given the formula $\Omega(x,e^{-y},z,z^{\prime})=0$, we have the transformation formulas
@@
$\displaystyle\frac{\left(-1\right)^{n-1}}{\cos\left(\displaystyle\frac{1\pm i}{2}z\right)}\overset{z\rightarrow\left(1\mp i\right)\tilde{n}\pi}{=}\displaystyle\frac{1\mp i}{z-\left(1\mp i\right)\tilde{n}\pi}-2\displaystyle\sum_{k=1}^{\infty}{\displaystyle\frac{\overline{\zeta}\left(2k\right)}{\pi^{2k}}\left(\displaystyle\frac{1\pm i}{2}\right)^{2k-1}\left(z-\left(1\mp i\right)\tilde{n}\pi\right)^{2k-1}},$ (2.20)
where $\tilde{n}:=n-\displaystyle\frac{1}{2}$. Here $\overline{\zeta}(s)$ stands for the alternating Riemann zeta function, which is defined by
$\bar{\zeta}\left(s\right):=\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{\left(-1\right)^{n-1}}{n^{s}}}\quad\left(\Re\left(s\right)>0\right).$
@@
([25, Lem. 1]) Let $p$ be a positive integer and $\alpha,\beta$ be real numbers such as $\alpha\beta=\pi^{2}$
@@
Lemma 2.6.
For any positive even integer $k$, we have
$\left(\sqrt{2}+1\right)^{k}+\left(\sqrt{2}-1\right)^{k}\in\mathbb{Z},\quad\sqrt{2}\left[\left(\sqrt{2}+1\right)^{k}-\left(\sqrt{2}-1\right)^{k}\right]\in\mathbb{Z},$ (2.24)
and for any positive odd integer $k$, we have
$\left(\sqrt{2}+1\right)^{k}-\left(\sqrt{2}-1\right)^{k}\in\mathbb{Z},\quad\sqrt{2}\left[\left(\sqrt{2}+1\right)^{k}+\left(\sqrt{2}-1\right)^{k}\right]\in\mathbb{Z}.$ (2.25)
Proof.
Expanding the expressions in (2.24) and (2.25) via the binomial theorem proves the lemma. ∎
@@ 3 Mixed Berndt-Type Integrals via Hyperbolic Series
In this section, we establish some explicit relations between mixed Berndt-type integrals and some hyperbolic summations using the contour integrations.
@@
where $C_{R}$ denotes the quarter- circular contour consisting of the interval $[0,R]$, the quarter-circle $\Gamma_{R}$ with $|z|=R,0\leq\arg(z)\leq\pi/2$, and $[iR,0]$, see the following figure:
Obviously, there exist poles when
$\left(\sinh z+i\sin z\right)\left(\cosh z+\cos z\right)=2i\cos^{2}\left\{\displaystyle\frac{1+i}{2}z\right\}\sin\left\{\left(1-i\right)z\right\}=0.$
The only poles lying inside $C_{R}$ are $z_{n}=\displaystyle\frac{n(1+i)\pi}{2}$ for $n\geq 1$ with $|z_{n}|<R$. Using Lemma 2.3, the residue $\text{Res}[F_{1}(z),z_{n}]$ at each such pole is given by
@@
As $R\rightarrow\infty$, it is easy to show that
@@
Thus, by combining the related identities, we deduce the desired result. ∎
Remark 3.2.
Fix the principal branch of $z^{a}$ on the sector $0<\arg z<\pi/2$. For $0<\Re(a)<1$, indent the contour at the origin by the small arc $\Gamma_{\varepsilon}=\{\varepsilon e^{it}:0\leq t\leq\pi/2\}$, see the following figure. Near $z=0$ one has $F_{1}(z)=O(|z|^{\Re(a)-1})$, hence
$\Bigl|\displaystyle\int_{\Gamma_{\varepsilon}}F_{1}(z)\,dz\Bigr|=O(\varepsilon^{\Re(a)})\to 0,\quad\text{as}\,\varepsilon\to 0.$
The large circular arc still gives $o(1)$. Therefore the left-hand side of (3.1) defines an analytic function of $a$ on $\Re(a)>0$. The right-hand side is entire in $a$ (uniform convergence). Since the identity holds for $a\geq 1$, the identity theorem extends (3.1) to the half-plane $\Re(a)>0$. Likewise, by the same argument, (3.6) extends to $\Re(a)>0$.
@@
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{a}\left(\sinh x-i\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}+$ $\left(-i\right)^{a}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{a}\left(\sin x-i\sinh x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$
@@
where $C_{R}$ denotes the quarter-circular contour consisting of the interval $[0,R]$, the quarter-circle $\Gamma_{R}$ with $|z|=R\ (-\pi/2\leq\arg(z)\leq 0)$, and [iR,0][-iR,0], see the following figure.
Similar to Theorem 3.1, we derive that the only poles inside $C_{R}$ are $z_{n}=\displaystyle\frac{n(1-i)\pi}{2}$ for $n\geq 1$ with $|z_{n}|<R$, and the residue $\text{Res}[F_{2}(z),z_{n}]$ at each pole is given by
$\mathrm{Res}\left[F_{2}\left(z\right),z=\displaystyle\frac{1-i}{2}n\pi\right]=\left(\displaystyle\frac{1-i}{2}\right)^{a+1}\displaystyle\frac{\pi^{a}}{2i}\displaystyle\frac{\left(-1\right)^{n-1}n^{a}}{\sinh^{2}\left(\displaystyle\frac{n\pi}{2}\right)}.$ (3.31)
It is easy to show that, letting $R\rightarrow\infty$,
@@
$-2\pi i\displaystyle\sum_{n=1}^{\infty}{\mathrm{Res}\left[F_{2}\left(z\right),z_{n}\right]}=\displaystyle\lim_{R\rightarrow\infty}\displaystyle\int_{\mathrm{C}_{R}}{\displaystyle\frac{z^{a}\mathrm{d}z}{\left(\sinh z+i\sin z\right)\left(\cosh z-\cos z\right)}}$
$=\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{a}\left(\sinh x-i\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}+\displaystyle\int_{0}^{\infty}{\displaystyle\frac{(-i)^{a}x^{a}\left(\sin x-i\sinh x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}.$
Thus, the proof is complete. ∎
Remark 3.4.
Fix the principal branch of $z^{a}$ on the sector $\pi/2<\arg z<0$. For $2<\Re(a)<3$, indent the contour at the origin by the small arc $\Gamma_{\varepsilon}=\{\varepsilon e^{it}:-\pi/2\leq t\leq 0\}$, see the following figure. Near $z=0$ the integrand in (3.3) satisfies $F_{2}(z)=O(|z|^{\Re(a)-3})$, hence
$\Bigl|\displaystyle\int_{\Gamma_{\varepsilon}}F_{2}(z)\,dz\Bigr|=O(\varepsilon^{\Re(a)-2})\to 0,\quad\text{as}\,\varepsilon\to 0,$
and the large arc still contributes $o(1)$. Therefore the left-hand side of (3.3) defines an analytic function of $a$ on $\Re(a)>2$. The right-hand side is entire in $a$ (uniform convergence). Since the identity holds for all real $a\geq 3$, the identity theorem extends (3.3) to the half-plane $\Re(a)>2$. The case of (3.5) follows by the same argument.
@@
where $C_{R}$ denotes the same positively oriented quarter-circular contour that we used in the proof of Theorem 3.3. The proof is completely similar as the proof of Theorem 3.3, so we leave it to the interested reader. ∎
@@
where $C_{R}$ denotes the same positively oriented quarter-circular contour that we used in the proof of Theorem 3.1. The proof is completely similar as the proof of Theorem 3.1, we leave the details to the interested reader. ∎
Setting $a=2\mu+1$ and $2\mu$ (or alternatively $2\mu+2$) in Theorems 3.1-3.6 respectively, yields the following corollaries.